The perimeter of a rectangle is 38 inches. If the length is 3 inches more then the witdth, find the width?
Answer by checkley77(12844)

L=W+3
2L+2W=38
2(W+3)+2W=38
2W+6+2W=38
4W=38-6
4W=32
W=32/4
W=8 ANS. FOR THE WIDTH.
L=8+3
L=11 ANS. FOR THE LENGTH.
PROOF:
2*11+2*8=38
22+16=38
38=38